1.In a bag of marbles you have 6 green, 4 yellow, 8 blue, and 2 red marbles. What is the probability of picking a yellow marble from the bag?

2.Theoretical probability is approximated by conducting trials and recording the ration of the number of occurrences o the event to the number of trials. true or false?

3.What is the probability of rolling a 2 or a 3 with a six sided die
Answer #1

1:

Total number of marbles = 6+4+8+2 = 20

Out of 20 marbles, 4 are yellow so he probability of picking a yellow marble from the bag is

P(yellow) = 4 /20 = 0.20

2:

Theoretical probability is probability of ration of number of favorable outcomes to all possible outcomes.

It is false.

3:

When we roll a six sided die possible outcomes are 1, 2, 3, 4, 5, and 6. Out of 6 outcomes, 2 outcomes are 2 or 3. So

P(2 or 3) = 2/ 6= 1/3